# Continual Learning with Query-Only Attention

### v1

@@ Abstract
Continual learning involves learning from a stream of data without repetition of data points, a scenario that is inherently complex due to distributional shift across tasks. We propose a query-only attention mechanism that discards keys and values, yet preserves the core inductive bias of transformer architectures. In continual learning scenarios, this simplified mechanism significantly mitigates both loss of plasticity and catastrophic forgetting, outperforming baselines such as selective re-initialization. We establish a conceptual link between query-only attention, full transformer attention, and model agnostic meta-learning, framing them as instances of meta-learning. We further provide intuition for why query-based models and attention networks help preserve plasticity in continual settings. Finally, through preliminary Hessian spectrum analysis, we observe that models maintaining higher curvature rank across tasks tend to retain plasticity. Our findings suggest that full attention may not be essential for capturing the benefits of meta-learning in continual learning.
@@ 1 Introduction
  • We provide a conceptual explanation of why the query-based model mitigates both loss of plasticity and catastrophic forgetting by converging to a global solution, rather than task-specific local solutions.
  • We show how the query-based model relates to the full attention mechanism used in transformers.
  • We establish a connection between our approach and meta-learning algorithms [7], in particular showing how it adapts ideas from MAML [6] to continual learning.
  • We analyze Hessian spectra and effective rank [13], demonstrating that if rank decreases across tasks correlates with loss of plasticity.
  • We demonstrate superior performance over multiple baselines in mitigating loss of plasticity.
  • Our results further show that catastrophic forgetting is reduced naturally by the query-based model, although eliminating forgetting is not the primary focus of this work.
@@ 2 Related work
Hence, continual learning faces challenge on two fronts, forward performance / mitigating loss of plasticity and backward performance / mitigating catastrophic forgetting. [2] showed that there exists a tradeoff between the two. We show an alternative view that query_based_cl and related architectures attain a global solution which mitigates both loss of plasticity and catastrophic forgetting simultaneously.
@@ 4 Problem statement
During inference, when the model is at data point $m$ of task $t$, the goal is to predict the next $n$ points $\{m+1,\dots,m+n\}$ from the same task $t$. The accuracy (or loss) on these future points defines the forward performance. A degradation in forward performance over time is taken as evidence of loss of plasticity.
For backward performance, after the model has been trained up to task $t$, we evaluate it on a previous task tjt-j by sampling $n$ data points from task tjt-j. The degradation in accuracy (or increase in loss) on these past tasks quantifies the extent of catastrophic forgetting.
  • Forward testing (plasticity): The model is evaluated using support examples already present in the replay buffer. No task identifiers are needed, since the buffer itself provides sufficient context for adaptation.
  • Backward testing (forgetting): The model is evaluated on a small held-out set of examples from past tasks. Here task identities are assumed to be known, so that performance on specific earlier tasks can be measured.
During testing:
  • Forward testing (plasticity): When the model is at data point $m$ of task $t$, it is evaluated on the next $n$ points $\{m+1,\dots,m+n\}$ from the same task $t$. Task identities are not available in this evaluation, so the model must rely only on the stream itself. Degradation in this metric across tasks indicates loss of plasticity.
  • Backward testing (forgetting): After the model has been trained up to task $t$, it is evaluated on a previous task tjt-j by sampling $n$ data points from $\mathcal{D}_{t-j}$. Here we assume access to a small buffer of examples from past tasks, and thus task identifiers are known at evaluation time. Degradation in performance on these past tasks quantifies catastrophic forgetting.
@@ 5.2 MAML with replay buffer
An important property of this setup is that the support/query sizes remain much smaller than those required by attention-based or query-only models. This makes MAML particularly well-suited for backward evaluation (catastrophic forgetting), since only a small set of examples per task is needed to adapt. Standard MAML inner-loop adaptation and outer-loop meta-updates are then applied on these partitioned tasks.
@@
1:Stream $(x_{t},y_{t})$; replay buffer $\mathcal{B}$ of size $N$; tasks $t$; support $s$; query $q$; learning rates $\alpha,\beta$
@@ 6.1 Global vs local solution
Comparing Equation 7 with Equation 6, the difference lies in how the distance metric is obtained: in kkNN it is fixed manually, while in the query-based model it is learned as $\theta$. Once $\theta$ is learned, predictions depend primarily on the support set, so continual adaptation occurs in-context rather than through constant parameter updates. This makes the learning global in nature and independent of any single task which is different from vanilla backpropagation or continual backpropagation algorithm where the model updates its parameters continuously local for each task.
@@ 6.2 Model agnostic meta-learning for continual learning
From Equation 10, predictions depend on task-specific parameters $\theta^{\prime}$ generated at inference time. This parallels the task-specific adaptation in MAML’s inner/outer-loop updates (Equations 3, 4). However, unlike query-based models, MAML was originally designed with task IDs and distributions known, an assumption that breaks in continual learning. In our experiments, we find that using a large replay buffer stabilizes MAML training despite this limitation. Even more interestingly, MAML requires a much smaller support set than query-based or attention networks, which could make it a promising direction for mitigating catastrophic forgetting in future continual learning work where memory efficiency is crucial.
@@ 7 Experiments
Figure 1: Permuted MNIST results (mean over 3 seeds; shaded regions indicate $\pm 1$ standard deviation across seeds).
@@
Figure 2: SCR results (mean over 3 seeds; shaded regions indicate $\pm 1$ standard deviation across seeds).
We evaluate forward (plasticity) and backward (forgetting) performance on two benchmarks: Permuted MNIST (abrupt distribution shifts) and Slowly Changing Regression (SCR) (gradual drift). All results are averaged over 3 seeds; shaded regions show $\pm$1 std. Higher accuracy (classification) / lower MSE (regression) is better.
Baselines and support sizes.
We compare our query-based model to vanilla backpropagation (BP), continual backpropagation (CBP), and attention network. On Permuted MNIST, we report two query-based variants: V1 with support size $1000$ and V2 with support size $200$. The transformer uses support size $100$; larger sets are prohibitive due to its $\mathcal{O}(n^{2})$ attention, whereas our query-only method is $\mathcal{O}(n)$. We also include a MAML-style model with a large replay buffer and a small support set of $10$ (trained on $100$ tasks for cost). On SCR, we use a single query-based model with support size $100$, matching the transformer (both $100$) to isolate algorithmic effects under equal memory/compute.
Observations on Permuted MNIST.
Query-based (V1/V2) achieves the highest forward accuracy and the lowest forgetting, and converges faster than the transformer. The transformer reaches comparable final accuracy but is slower due to quadratic cost. BP shows gradual loss of plasticity; CBP improves stability but still lags. MAML converges quickly with strong accuracy despite being trained on only $100$ tasks.
Observations on SCR.
With equal support ($100$) for query-based and transformer, both sustain low MSE in forward testing. BP gradually loses plasticity, and CBP does not learn well on this benchmark. In backward testing, BP/CBP exhibit severe catastrophic forgetting, while query-based and MAML retain performance.
Note that here, including $y_{i}$ in $Q_{\theta}(x_{t},x_{i},y_{i})$ did not significantly improve performance, so we used only $(x_{t},x_{i})$ to reduce computation. In contrast, for Permuted-MNIST the inclusion of labels $y_{i}$ in $Q_{\theta}$ yielded consistently better performance, and was therefore retained.
@@
Across both benchmarks, query-based and full attention behave similarly, but query-based converges faster and often performs better because it leverages larger support in $\mathcal{O}(n)$ time (Permuted MNIST), and remains competitive even under equal support (SCR). This highlights practical advantages for continual learning where compute and memory are tight.
@@ 7.1 Effective rank
Figure 3: Permuted-MNIST. Mean $\pm$ std over 3 seeds. Effective rank co-varies with forward performance; dips align with reduced plasticity.
The effective rank serves as a proxy for plasticity. We measure it at the start of each task using the Hessian of the final layer only, since full-Hessian computation is infeasible. On Permuted-MNIST, vanilla backpropagation shows a steady drop in effective rank, aligning with loss of plasticity. Other models maintain more stable ranks, indicating preserved plasticity. Note that absolute rank values differ across models due to variation in the size of their final layers, which is the only layer we analyze. Due to computational constraints, this analysis was not extended to SCR.
@@ 8 Conclusion
We introduced a query-only attention mechanism for continual learning, showing that it mitigates both loss of plasticity and catastrophic forgetting while reducing computational cost compared to full attention. Our analysis connects query-only models to MAML and full attention through the lens of global vs. local solutions, and further relates them to $k$-nearest neighbors. Hessian rank experiments support the role of curvature in sustaining plasticity. While our evaluation is limited to permuted MNIST and regression benchmarks, the results are promising.
@@ A.3 Permuted MNIST Setup
Table 1: Permuted MNIST configuration: Query-based CL (V1).
@@
Table 2: Permuted MNIST configuration: Query-based CL (V2).
@@
Wall-clock 0.5 hours (ran only first 100 tasks)
@@
A.4 Slowly changing regression (SCR)
@@
Table 7: Slowly changing regression task: Query-based CL.
@@
Wall-clock 1 hour to run 100 tasks

### v2

@@ Abstract
Continual learning involves learning from a stream of data without repetition of data points, a scenario that is inherently complex due to distributional shift across tasks. We propose a query-only attention mechanism that discards keys and values, yet preserves the core inductive bias of full-attention architectures. In continual learning scenarios, this simplified mechanism significantly mitigates both loss of plasticity and catastrophic forgetting, outperforming baselines such as selective re-initialization. We establish a conceptual link between query-only attention, full transformer attention, and model agnostic meta-learning, framing them as instances of meta-learning. We further provide intuition for why query-only attention and full-attention networks help preserve plasticity in continual settings. Finally, through preliminary Hessian spectrum analysis, we observe that models maintaining higher curvature rank across tasks tend to retain plasticity. Our findings suggest that full attention may not be essential for capturing the benefits of meta-learning in continual learning.
@@ 1 Introduction
  • We introduce Query-Only Attention, which is based on attention mechanism and meta-learning that mitigates loss of plasticity more effectively than state-of-the-art methods in fully online continual learning experiments.
  • As a natural consequence, Query-Only Attention also mitigates catastrophic forgetting when task identity is available, although forgetting reduction is not the primary focus of this work.
  • We provide a conceptual explanation showing that Query-Only Attention mitigates both loss of plasticity and forgetting by converging toward a global rather than task-specific local solution.
  • We establish conceptual links between Query-Only Attention, the original attention mechanism [21], and meta-learning approaches such as MAML [6] in continual learning context.
  • We analyze the Hessian spectrum and effective rank [12], demonstrating that decreasing rank across tasks correlates with loss of plasticity.
@@ 2 Related work
Hence, continual learning faces challenge on two fronts, forward performance / mitigating loss of plasticity and backward performance / mitigating catastrophic forgetting. [2] showed that there exists a tradeoff between the two. We show an alternative view that Query-Only Attention and related architectures attain a global solution which mitigates both loss of plasticity and catastrophic forgetting simultaneously.
@@ 4 Problem statement
Evaluation Protocol. During inference, the model processes a data stream sequentially. At data point $m$ of task $t$, the goal is to predict the next $n$ points $\{m+1,\dots,m+n\}$ from the same task. The resulting accuracy (or loss) defines the forward performance; its degradation over time indicates loss of plasticity.
After training up to task $t$, the model is also evaluated on samples from a previous task tjt-j. The resulting accuracy defines the backward performance, and its degradation quantifies catastrophic forgetting.
  • Forward testing (plasticity): Evaluates on upcoming data from the current stream without task identifiers. A decline in this metric across tasks signals loss of plasticity.
  • Backward testing (forgetting): Evaluates on a small held-out buffer of past-task samples where task identities are known. A decline in this metric indicates catastrophic forgetting.
@@ 5.2 MAML with replay buffer
An important property of this setup is that the support/query sizes remain much smaller than those required by attention-based or query-only models. This makes MAML particularly well-suited for backward evaluation (catastrophic forgetting), since only a small set of examples per task is needed to adapt. Standard MAML inner-loop adaptation and outer-loop meta-updates are then applied on these partitioned tasks. The details of the algorithm are presented in the appendix section.
@@
1:Stream $(x_{t},y_{t})$; replay buffer $\mathcal{B}$ of size NN; tasks $t$; support $s$; query $q$; learning rates $\alpha,\beta$
@@ 6.1 Global vs local solution
Comparing Equation 7 with Equation 6, the difference lies in how the distance metric is obtained: in kkNN it is fixed manually, while in the query-only attention it is learned as $\theta$. Once $\theta$ is learned, predictions depend primarily on the support set, so continual adaptation occurs in-context rather than through constant parameter updates. This makes the learning global in nature and independent of any single task which is different from vanilla backpropagation or continual backpropagation algorithm where the model updates its parameters continuously local for each task.
@@ 6.2 Model agnostic meta-learning for continual learning
From Equation 10, predictions depend on task-specific parameters $\theta^{\prime}$ generated at inference time. This parallels the task-specific adaptation in MAML’s inner/outer-loop updates (Equations 3, 4). However, unlike query-only attention models, MAML was originally designed with task IDs and distributions known, an assumption that breaks in continual learning. In our experiments, we find that using a large replay buffer stabilizes MAML training despite this limitation. Even more interestingly, MAML requires a much smaller support set than query-only attention or full attention networks, which could make it a promising direction for mitigating catastrophic forgetting in future continual learning work where memory efficiency is crucial.
@@ 7 Experiments
We evaluate forward (plasticity) and backward (forgetting) performance on three benchmarks: Permuted MNIST (abrupt shifts), Tiny ImageNet, and Slowly Changing Regression (SCR) (gradual drift). The above experiments follow same setup as explained in Dohare et al. [3], except instead of full ImageNet we choose Tiny ImageNet for efficiency. Distinction is that we perform experiments in a online setting, using unknown task boundaries to study loss of plasticity and known task boundaries to study catastrophic forgetting. Detailed configurations appear in the Appendix. Results are averaged over three seeds with shaded $\pm$1 std regions. Higher accuracy (classification) and lower MSE (regression) indicate better performance. Baselines include BP, CBP, and the Full-Attention Network. Since the primary focus of this work is mitigating loss of plasticity, we additionally include the state-of-the-art forgetting baseline Elastic Weight Consolidation (EWC) in the ImageNet experiments for completeness.
7.1 Permuted MNIST
We evaluate the query-only attention model with support sizes of $1000$ and $200$ and name them as Query-Only Attention V1 and Query-Only Attention V2. The full-attention uses support $100$, since its $\mathcal{O}(n^{2})$ attention limits scalability, while our $\mathcal{O}(n)$ query-only design allows larger supports at lower cost. The MAML-style model uses a large replay buffer with a small support of $10$, trained on $100$ tasks. Each iteration is costlier due to inner-loop updates, so it runs on fewer tasks but performs multiple updates per iteration. For readability, performance curves are averaged over fixed task windows; results without averaging are included in the Appendix.
@@ 7.1 Permuted MNIST
Figure 1: The prediction is over 7500 tasks and each data-point in the graph is averaged over 100 tasks for all models except for MAML. For MAML, we run over only 75 tasks and is shown without averaging.
7.2 Split Image Net
Split-image-net we use a support size of 180 for both query-only attention and full-attention model. For MAML a support size of only size 10 is enough.
(a) Forward accuracy (Split Image Net).
(b) Backward accuracy (Split Image Net).
Figure 2: The prediction is over 9000 tasks and each data-point in the graph is averaged over 100 tasks for all the models except MAML. MAML is run over 500 tasks, averaged over 5 tasks.
Observations in Classification Tasks.
Across both 7.1 and 7.2, the query-only attention model consistently outperforms all baselines in forward and backward performance. The full-attention reaches similar final accuracy but with 50% more parameters and slower convergence. Under unknown task boundaries, CBP performs poorly, especially on Split ImageNet, while the vanilla network shows clear loss of plasticity. The MAML-based model converges fastest, achieving intermediate performance between attention-based and standard networks.
7.3 Slowly Changing Regression
In SCR, we use a single query-only attention model with support size $100$, matching the full-attention (both $100$) to isolate algorithmic effects under equal memory/compute. With equal support ($100$) for query-only attention and full attention, both sustain low MSE in forward testing.
@@ 7.3 Slowly Changing Regression
Figure 3: The prediction is over 800 tasks and each data-point in the graph is averaged over 10 tasks for all models.
Observations on Regression Task.
As in classification, BP gradually loses plasticity, while CBP fails to learn effectively due to its purely online setup. The query-only attention model converges quickly with near-zero MSE, whereas full attention converges more slowly with slightly lower performance. Unlike in classification, the MAML-based model struggles to converge but still maintains plasticity. In backward testing, BP and CBP show severe forgetting, while query-only attention, full-attention, and MAML models retain high performance. Including $y_{i}$ in $Q_{\theta}(x_{t},x_{i},y_{i})$ offered no gain on this regression task, so we used only $(x_{t},x_{i})$; for Permuted-MNIST, label inclusion improved results and was retained.
@@
Across all benchmarks, query-only attention and full-attention models perform similarly, but the query-only attention model converges faster and scales better with $\mathcal{O}(n)$ complexity. The MAML-based approach shows strong backward performance and quick convergence with minimal support, though less consistent on SCR. Query-only attention model mitigates loss of plasticity and forgetting, highlighting its practicality under limited compute and memory.
@@ 7.4 Effective Rank
Figure 4: Effective rank co-varies with forward performance; dips align with reduced plasticity. Mean $\pm$ std over 3 seeds.
The effective rank serves as a proxy for plasticity. We measure it at the start of each task using the Hessian of the final layer only, since full-Hessian computation is infeasible. We measure it only on Permuted MNIST and SCR and not on image-net due to computational constraints. Also, the effective rank has been normalized since effective rank will depend on size of neural net. In both Permuted-MNIST and SCR, vanilla backpropagation shows a steady drop in effective rank, aligning with loss of plasticity. All other models show near minimal drop in effective rank thus indicating preserved plasticity.
@@ 8 Conclusion
We introduced a query-only attention mechanism for continual learning, showing that it mitigates both loss of plasticity when task boundary is unknown beating state of the art models and no task repetition. If the task boundary is known, then query-only attention can also mitigate catastrophic forgetting. Query-Only Attention has lower computational cost compared to full attention. Our analysis connects query-only models to MAML and full attention through the lens of global vs. local solutions, and further relates them to $k$-nearest neighbors. We further confirmed this relationship by running on three experiments.
Hessian rank experiments support the role of curvature in sustaining plasticity across different approaches which mitigate loss of plasticity.
@@ A.3 Permuted MNIST Setup
Table 1: Permuted MNIST configuration: Query-Only Attention (V1).
@@
Table 2: Permuted MNIST configuration: Query-Only Attention (V2).
@@
Wall-clock 4.2 hours
@@
A.4 Split Image Net
We adopt the Split Image Net benchmark introduced by Dohare et al. [3]. However, due to computational constraints we utilize tiny version of the image net instead of full image net. We further downsample the images to 32 * 32 for faster computation. The remaining setup is same as in the original paper, ensuring comparability of results. For full details, we refer readers to the experimental protocol in [3]. Tiny image net consits of 200 labels and 500 images per label in training setup. The test setup consists of 50 labels per class. To incorporate tiny image net for continual learning, each task consits of randomly sampling images from 2 classes. Thus a task is a binary classification task with 1000 datapoints for training. At the end of training in that class we measure the accuray on 100 datapoints corresponding to 2 lables used in training. Since the training is purely online fashion a task is trained for a single epoch and then validation is carried out, followed by images for next task. All the below architectures presented use the same CNN architure as the starting point for input image transformation:
Table 7: Split Image Net: CNN architecture (same for all models)
Setting Value
(Input channels, Output Channels,Kernel Size, Padding, MaxPool, Activation) ( 3, 32, 5, 1, 2, Relu )
(Input channels, Output Channels,Kernel Size, Padding, MaxPool, Activation) ( 32, 64, 3, 1, 2, Relu )
(Input channels, Output Channels,Kernel Size, Padding, MaxPool, Activation) ( 64, 128, 3, 1, 2, Relu )
Init Xavier uniform (weights), Zeros (biases)
Table 8: Split Image Net: Query-Only Attention.
Setting Value
Support size ($|B|$) 180
Replay buffer size 200
Distance metric Learned $Q_{\theta}$ (query-only)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching (batch size =10=10)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 9500
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) ( 2304, 128, 1, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 3 hours/run
Table 9: Split Image Net: MAML.
Setting Value
(Support size, query size, tasks per iteration) ($|B|$) (10, 10, 5)
Replay buffer size 20000
Optimizer Adam
Outer Learning rate 1e-4
Inner Learning rate 1e-2
Weight decay 0.0
Batching batch size =10=10
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 500
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 1152, 128, 2, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 5 hours (ran only first 100 tasks)
Table 10: Split Image Net: Attention Network baseline.
Setting Value
Support size ($|B|$) 180
Attention Full self-attention ($\mathcal{O}(n^{2})$)
Replay buffer size 200
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching batch size =10=10
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 9500
Steps per task 150
Architecture 3 layers, 1 head, d_model=130
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 4.5 hours/run
Table 11: Split Image Net: Vanilla Backpropagation.
Setting Value
Support size N/A
Replay buffer size 0 (no replay)
Optimizer Adam
Learning rate 1e-4
Weight decay 0.0
Batching Online (batch size =10=10)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 9500
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 1152, 128, 2, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 4.0 hours/run
Table 12: Split Image Net: Continuous Backpropagation.
Setting Value
Support size N/A
Replay buffer size 0 (no replay)
Optimizer Adam
Learning rate 1e-4
Weight decay 0
Batching Online (batch size =1=1)
Seeds 20, 30, 40 (report mean $\pm$ std)
Tasks 7000
Steps per task 150
(Input size, Hidden Size, Output Size, Hidden Layers) : ( 1152, 128, 2, 1 )
Init Xavier uniform (weights), Zeros (biases)
[…]
Wall-clock 5.5 hours/run
A.5 Slowly changing regression (SCR)
@@
Table 13: Slowly changing regression task: Query-Only Attention.
@@
Wall-clock 8.5 hour to run 100 tasks